OpenStudy (anonymous):

Solve the equation for x: 14x2 - 28x = 0 {-14, 28} {2, -7} {0, 2} no solution

OpenStudy (saifoo.khan):

C for Cat.

OpenStudy (anonymous):

zero is pretty clearly one solution because each term has an x in it

OpenStudy (sasogeek):

{0, 2}

OpenStudy (anonymous):

if you want to now if 2 works, check by replacing x by 2 you get \[14\times 2^2-28\times 2=56-56=0\]